Three times a number, decreased by 8 is the same as twice the number, increased by 15?Find the number.
Answer by zeynep(43)

Let's convert this word prolem into a mathematical expression:
Let the number be x.
Three times x is 3x.
If we decrease 3x by 8, it is 3x-8.
3x-8 is the same as twice the number, which is 2x, increased by 15, which is 2x+15.
So,
3x-8 = 2x+15
To solve the equation, move 2x to the left of the equation and -8 to the right of the equation.
3x-2x = 8+15
x = 23
The number is 23.